### v1

# Benchmarking Android Malware Detection: Rethinking the Role of Traditional and Deep Learning Models

Abstract
Android malware detection has been extensively studied using both traditional machine learning (ML) and deep learning (DL) approaches. While many state-of-the-art detection models, particularly those based on DL, claim superior performance, they often rely on limited comparisons, lacking comprehensive benchmarking against traditional ML models across diverse datasets. This raises concerns about the robustness of DL-based approaches’ performance and the potential oversight of simpler, more efficient ML models. In this paper, we conduct a systematic evaluation of Android malware detection models across four datasets: three recently published, publicly available datasets and a large-scale dataset we systematically collected. We implement a range of traditional ML models, including Random Forests (RF) and CatBoost, alongside advanced DL models such as Capsule Graph Neural Networks (CapsGNN), BERT-based models, and ExcelFormer-based models. Our results reveal that while advanced DL models can achieve strong performance, they are often compared against an insufficient number of traditional ML baselines. In many cases, simpler and more computationally efficient ML models achieve comparable or even superior performance. These findings highlight the need for rigorous benchmarking in Android malware detection research. We encourage future studies to conduct more comprehensive benchmarking comparisons between traditional and advanced models to ensure a more accurate assessment of detection capabilities. To facilitate further research, we provide access to our dataset, including app IDs, hash values, and labels.
Index Terms:
Android malware detection, Machine learning, Deep learning, Dataset, Benchmark
1 Introduction
The Android ecosystem maintains the leading position as the largest and most active system among the worldwide mobile computing environment by far. Among its ecosystem components, the app markets which consist of Google Play, as well as manufacturer and third-party app stores, play a crucial role in releasing and distributing the apps to end users. According to one recent 42matters’s statistics [1], there were 40,325 apps on average released monthly just through Google Play Store between November 2024 and Juanuary 2025. The app market’s rapid growth makes it crucial to have an automated and effective app vetting system that scales to the app market sizes. An important component of a vetting system is the application of machine learning to triage the apps in the vetting process.
Traditional machine learning approaches (shorthanded ML hereafter) such as Naive Bayes (NB), k-Nearest Neighbors (k-NN), Support Vector Machines (SVM), Random Forest (RF), Artificial Neural Networks (ANN), Gradient Boosting, etc. have been applied to train models on a training dataset, which consists of apps whose maliciousness is known. Then the trained models are applied to classify a test dataset, where apps’ maliciousness is unknown. Many traditional ML models have shown outstanding performance [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] on datasets from small to market-scale.
More recently, deep learning approaches (shorthanded DL hereafter) have been studied on Android malware detection [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. One potential advantage of DL over ML is that it does not require an extensive feature selection and engineering process, although careful consideration still needs to be applied as to how to represent app data provided as input to the DL models.
Despite the proposed DL approaches for this problem, there exist limited comparisons [12, 21, 27, 24, 26] between ML and the proposed DL models. More often than not researchers do not publish the dataset used to compare performances of the various models, making it impossible for other researchers to further benchmark the model against additional alternatives. It is encouraging that we are seeing some recent works where researchers publish the datasets used in their study. Upon looking through the literature, we identified two works where 1) the researchers compared novel DL models against ML models when applied to Android malware detection, 2) the researchers published the datasets used in their experiments, and 3) the datasets are of substantial sizes. This enabled us to conduct additional experiments to benchmark the proposed DL models against additional ML models that were not examined in the original research. In addition, we conducted benchmarking of two other DL models against the ML models, on a large dataset that we curated. Our experiments show mixed results regarding whether DL models can outperform ML models for Android malware detection. The results call for more extensive benchmarking of new deep learning models against traditional machine learning models for Android malware detection.
2 Related Work
Machine learning-based Android malware detection systems have been researched for a decade. Crowdroid [2] is an earlier ML framework for Android malware detection. DroidAPIMiner [3] performs static analysis to extract lightweight features at API level, then uses different classification models such as k-NN, SVM and Decision Trees (DT) to detect malware. DREBIN [6] uses similar static analysis techniques as DroidAPIMiner to generate a feature set that contains components, permissions, filtered intents, API calls and network addresses, then uses SVM as a malware classifier. Roy et al. [28] performed experimentation studies to investigate a number of research questions related to using ML in Android malware detection. MaMaDroid [29] abstracts apps’ API calls from 340 packages and 11 families of malware, then applies a Markov chain to form the sequence of abstracted API calls extracted from call graphs as a feature vector, and finally performs classification using Random Forest, k-NN and SVM algorithms. APIChecker [11] uses dynamic analysis to collect API features, applies nine machine learning algorithms for Android malware detection and evaluates their performance on the T-Market data with about 500K apps. The results show that RF yields the best overall performance among the selected models. DroidAPIMiner [4], DREBIN [6], DroidMat [30], and MaMaDroid [29] also compared their performance with other traditional ML classifiers and anti-virus scanners. MaMaDroid uses four different ML algorithms and the detection results obtained with RF have better performance in general.
Recent years have seen an increase in deep learning related Android malware detection research. The deep neural network models used include deep belief networks (DBN) [12, 13], convolutional neural networks (CNN) [15, 16, 24, 31], pre-trained BERT models [26], recurrent neural networks (mainly long short-term memory networks, or LSTM) [32, 19], and multilayer perceptrons (MLP) with continuous learning schemes [25], among others. Droid-Sec [12] trained and tested a DBN model based on 500 apps (250 malware). Experimental results showed that the DL model has significant accuracy improvement as compared to traditional ML models, such as SVM, NB, Logistic Regression, and MLP. Deep4MalDroid [14] evaluated a Stacked AutoEncoders (SAEs) model based on 3000 apps (50% of which are malware). Results showed that the proposed DL model has higher accuracy than the selected ML models, specifically, SVM, ANN, NB, and DT. DeepFlow [33] proposed a DBN model, and conducted experiments based on 3000 benign apps and 8000 malicious apps. The comparison between DL and traditional ML models showed that DBN and SVM outperform other selected ML models in balancing the recall and precision. DexRay [24] converts the raw bytecode of DEX files in the APK as one stream vector, then converts the vector to a grey-scale image which is provided as input to a CNN architecture. DexRay evaluated the performance of the CNN model based on 61,809 malicious and 134,134 benign apps, then compared its detection results with DREBIN’s results and other image-based malware detection approaches. DexRay yields almost the same performance metrics as DREBIN; however, its performance can be significantly affected if obfuscated apps are absent in the training data but present in the test data. DetectBERT [26] leverages a pre-trained BERT-like model with correlated Multiple Instance Learning (c-MIL) to process Smali code from APKs, enriching class-level features and aggregating them into an app-level representation for effective Android malware detection. It employs the DexRay dataset as its benchmark baseline for evaluation. DeepRefiner [17] is a two-layer detection system, with the first detection layer based on MLP, and the second layer based on an LSTM network. The experimental data include 47,525 benign apps and 62,915 malicious apps. The detection performance, based on accuracy, TPR and FPR, is approximately 10% higher than that of a fine-turned SVM model. DANdroid [34] uses the DREBIN dataset to evaluate its proposed DL discriminative adversarial network, and the performance is very close to that of DREBIN. DL-Droid [23] generates 420 dynamic stateful features such as application attributes, actions and permissions by running apps on real devices, then feeding them into a multilayer perceptron classifier. It uses weighted F measure (w-FM) as main evaluation metric and compares it against that of seven selected ML algorithms. DL-Droid reached the highest w-FM value of 98.82% (the best ML model RF reached w-FM value of 97.1%). AMD-CNN [31] is a visualization-based Android malware detection tool. It extracts the features from Android manifest file and converts them to RGC images, then feeds the images to a CNN network. It was evaluated based on 960 benign and 960 malicious apps. [35] identified five factors that influence the training and performance of ML-based Android malware detectors. The authors conducted a fair comparison of eight detectors on a mixed-market dataset, accounting for these factors. The detectors include seven traditional ML models (across four types) and one DBN with two hidden layers of 150 neurons each. Chen et al. [25] proposed a continuous learning approach using MLP to enhance the adaptability of ML-based Android malware detectors to concept drift [36]. Their work includes a comprehensive comparison with major baseline models, with both input data and model-related code available as open source. The evaluation was conducted using two established datasets, APIGRAPH [37] and AndroZoo [38].
Our study complements these prior works by extensive benchmarking the proposed DL models against additional standard ML models that were not examined by the original researchers. Our results lead to reassessment of DL’s advantage against ML in Android malware detection, and highlights the need for extensive benchmarking for this problem domain.
3 Benchmarking Methodology
Benchmarking across different models is a standard approach for evaluating and comparing their performance, as well as identifying the strengths and weaknesses of various approaches. However, determining the sufficient number of models to compare with is not easy, as further benchmarking using more models could always be beneficial. For this reason, it is utterly important that researchers make the dataset upon which their models are evaluated available. This will enable other researchers to evaluate additional models that had not been considered in the original researcher. More extensive evaluations can provide deeper insights and help understand the true strengths of the various models for Android malware detection.
One of the key challenges for such extensive evaluation is that not all works publish the datasets used in their study. Fortunately, we have identified recent studies, such as Chen et al.’s Enc+MLP [25] and Sun et al.’s DetectBERT [26], that have made their datasets publicly available. This allowed us conduct more extensive benchmarking of the models proposed in these two works. Additionally, we collected a Google Play-only dataset and used two DL approaches, Capsule Graph Neural Networks (CapsGNN) and ExcelFormer, which have not been explored before for Android malware detection. Each of the three works employs different datasets to evaluate their models, and their approaches to app representation also vary. Our primary objective is not to assert the superiority of any specific model but to demonstrate how comprehensive benchmarking enhances the understanding of model performance. To this end we adopt the following benchmarking strategy.
  1. 1.
    We keep the app representation (e.g., feature vectors) used in the original research. When the provided datasets include feature vectors used in learning, we simply use that information. When the original feature vectors are not publicly available, we generate them from the App apk files based on the information provided in the paper.
  2. 2.
    We use the exact same datasets as used in the original research, and benchmark additional ML models that were not examined in the original research.
Since each work uses different datasets, the second point indicates that when interpreting the results presented in this paper, comparisons shall only be made within each work. Comparison of performance results across different works is not meaningful. While it may be tempting to use the same dataset to evaluate all the models proposed in all these works, the research community has not agreed upon a “gold standard” dataset for Android malware detection. In fact to understand how choices in datasets construction impact models’ measured performance is a non-trivial research question in its own right and beyond the scope of this paper.
In the following sections, we first discuss traditional machine learning models considered and then explore additional models that were not included in the two prior studies.
4 Traditional Machine Learning Models Considered
Widely explored ML models in Android malware detection include Naive Bayes (NB) [39, 40], support vector machines (SVM) [4, 6], k-nearest neighbors (k-NN) [30, 40], and random forest (RF) [29]. We consider these four widely used models as potential baselines to compare with the proposed DL models. Additionally, Gradient Boosted Decision Trees (GBDTs) are powerful classification tools, with CatBoost [41] being a notable variant of GBDT-based ensemble techniques. Having been introduced in late 2018, CatBoost has demonstrated strong performance in classification and regression tasks across various domains, including cybersecurity [41]. Thus we also include it as one of the ML models in our benchmarking.
NB: This is a simple but effective classifier. Furthermore, it is the fastest classifier we evaluated. It can get the classification result quickly even with half a million apps.
SVM: Widely used classifier in Android malware classification. It requires more memory and computation time during the training stage, which poses a significant challenge for large datasets. We set the kernel to be linear, and use a grid search with values 0.001, 0.1, 1.0, and 10 to decide the optimal regularization parameter to train the SVM model.
k-NN: Another simple and robust traditional classifier. The value of the hyper-parameter $k$ has a direct impact on the performance. We use a grid search with values between 3 and 15 to determine the optimal $k$ value to train the k-NN model.
RF: A popular ensemble-type model based on decision trees. It has shown promising performance in Android malware detection [29] and has the capability to be applied to market-scale datasets [11]. We perform a grid search to find the optimal number of trees in the forest. The estimator value is between 100 and 300.
CatBoost: Another ensemble-based model built on decision trees. While it is relatively new compared to the other ML models, recent studies across various domains have demonstrated its effectiveness in classification tasks [41]. CatBoost is sensitive to hyperparameters, and our experimental results are based on the following setup: 1000 iterations, a learning rate of 0.1, a depth of 10, and an L2 regularization parameter (l2_leaf_reg) of 5.
5 Additional Benchmarking for DetectBERT
DetectBERT is a recent approach proposed by Sun et al. [26], which utilizes a pre-trained BERT-like model combined with Correlated Multiple Instance Learning (c-MIL) to process Smali code from APKs. This approach enhances class-level features and aggregates them into an app-level representation, improving the effectiveness of Android malware detection.
5.1 Benchmarking for DetectBERT
DetectBERT was evaluated against two state-of-the-art Android malware detection models, Drebin [6] and DexRAY [24]. Drebin utilizes SVM on static feature vectors, while DexRAY is a DL framework that transforms low-level bytecode into images and processes them using a CNN for malware detection. The results indicate that DetectBERT slightly outperforms both baseline models.
5.2 Dataset for DetectBERT
DetectBERT Benchmark Dataset. This large-scale dataset originates from the DexRay [24] study and comprises 96,994 benign apps and 61,809 malware apps. The labeling process is based on VirusTotal reports: apps that are not flagged by any antivirus engines are considered benign, while those detected as malicious by more than two antivirus engines are labeled as malware.
5.3 Additional Benchmarking
Model Accuracy Precision Recall F1 Score
Drebin 0.97 0.97 0.94 0.96
DexRay 0.97 0.97 0.95 0.96
DetectBERT 0.97 0.98 0.95 0.97
NB 0.85 0.87 0.85 0.86
RF 0.97 0.97 0.97 0.97
KNN 0.97 0.97 0.97 0.97
SVM 0.96 0.96 0.96 0.96
CatBoost 0.98 0.98 0.98 0.98
TABLE I: Performance comparison of various models on DetectBERT benchmark dataset. The unshaded are from the original paper, while the shaded results are from our benchmarking.
Sun et al. [26] open sourced their DetectBERT model for Android malware detection, which enables us to use the exact same dataset, data splits, and evaluation metrics for training and testing various models, ensuring fair comparisons.
One of the challenges in applying the DetectBERT model is generating the feature vector for each smali class in an APK file from a large-scale dataset. To address this, we explored the performance of more cost-effective approaches on this dataset. Since DREBIN is one of the baseline models used in the original research, we generated DREBIN-like static features for each app, including requested permissions, component names, intents, suspicious API calls, and network addresses. We then calculated the mutual information of the unique static features in this dataset and ranked them accordingly. By testing different feature vector sizes based on the top N features, we found that a feature vector consisting of 2,919 static features per app yielded optimal results. Table I presents the performance of various ML models against DetectBERT. The shaded rows represent the models we tested with these features. Except for NB, all other models (KNN, SVM, RF, and CatBoost) achieved performance comparable to DetectBERT. CatBoost slightly outperforms DetectBERT. These ML models also require significantly less compute time than DetectBERT.
In the paper the authors also conducted experiments on data sets that adhere to temporal consistency. However the published dataset does not include the train/test data split for this experiment, and as a result we were not able to compare additional ML models’ results against the results published in the paper.
6 Additional Benchmarking for Enc+MLP
Enc + MLP is proposed by Chen et al. [25] to address the concept drift [36] problem in Android malware detection. This scheme integrates contrastive learning with active learning, structuring the model into two subnetworks. The first subnetwork, a hierarchical contrastive encoder (Enc), employs contrastive learning techniques to encode input embeddings, ensuring that embeddings of the same malware family are closer to each other. The second subnetwork is an MLP classifier that utilizes these embeddings for malware classification.
The model is initially trained on a labeled dataset to establish the hierarchical contrastive classifier. Following this, an active learning process begins, where the trained classifier predicts labels for a batch of test samples. A pseudo-loss sample selector in the hierarchical contrastive classifier identifies the most uncertain apps within a predefined labeling budget and adds them to the training set for the next iteration. The model is then retrained using a warm start, and this process repeats continuously, refining the model with each subsequent batch of test data. This approach improves upon existing active learning baselines for Android malware detection while minimizing the need for manual labeling, as demonstrated through evaluations on two publicly available datasets.
6.1 Benchmarking for Enc+MLP
Enc+MLP uses multiple baseline models and explores uncertainty sampling for both binary and multiclass classifiers. The binary classifiers include a fully connected neural network (Binary MLP), a linear Support Vector Machine (Binary SVM), and Gradient Boosted Decision Trees (GBDT). The multiclass classifiers include an MLP and an SVM, and the authors also experimented with a combined classifier, “Multiclass MLP + Binary SVM.” Uncertainty is measured as one minus the maximum prediction score across all classes. All baseline models employ a cold-start approach (retraining from scratch on the updated training data) for active learning.
6.2 Datasets for Enc+MLP
Enc+MLP was evaluated on two datasets: APIGraph and AndroZoo.
Year
Malicious
Apps
Benign
Apps
Total
2012 3,061 27,472 30,533
2013 4,854 43,714 48,568
2014 5,809 52,676 58,485
2015 5,508 51,944 57,452
2016 5,324 50,712 56,036
2017 2,465 24,847 27,312
2018 3,783 38,146 41,929
TABLE II: APIGraph Dataset
Year
Malicious
Apps
Benign
Apps
Total
2019 4,542 40,947 45,489
2020 3,982 34,921 38,904
2021 1,676 13,985 15,662
TABLE III: AndroZoo Dataset
APIGraph Dataset. Collected by Chen et al. [25] using the app hash list provided by APIGraph [37], this dataset spans seven years of Android apps from 2012 to 2018. The apps are ordered based on their appearance timestamps in VirusTotal [42], addressing both spatial and temporal biases [43]. The dataset is evenly distributed across the years, with each year consisting of approximately 90% benign apps and 10% malicious apps. The malicious apps are sourced from VirusTotal [42], VirusShare [44], and the AMD dataset [45], while the benign apps are obtained from AndroZoo [38]. Table III provides a detailed breakdown of this dataset.
AndroZoo Dataset. Collected by Chen et al. [25] from AndroZoo [38], this dataset includes Android apps from 2019 to 2021.The apps are also ordered based on their appearance timestamps. Malicious apps were randomly selected based on VirusTotal reports, where at least 15 antivirus engines flagged them as malware. Benign apps were randomly chosen from those with 0 positive detections by any antivirus engine. The dataset maintains a similar malware/benign app ratio, with 90% benign apps and 10% malicious apps each year. Table III provides a detailed breakdown of the collected data for each year.
6.3 Additional Benchmarking
Monthly Sample Budget Model Architecture Sample Selector APIGraph Dataset Average Performance (%) AndroZoo Dataset Average Performance (%)
FNR FPR F1 FNR FPR F1
50 Binary MLP Uncertainty 23.77 0.52 83.84 53.12 0.46 59.50
Multiclass MLP Uncertainty 16.10 4.64 73.77 49.86 28.52 28.65
Multiclass MLP
+ Binary SVM
Uncertainty 38.40 1.01 71.38 73.13 2.87 34.04
Binary SVM Uncertainty 16.92 0.61 87.72 48.77 0.29 63.42
CADE OOD 36.11 12.9 71.70 62.01 0.55 50.26
multiclass SVM Uncertainty 35.79 0.17 87.43 65.77 0.09 46.91
multiclass GBDT Uncertainty 31.75 0.54 77.92 50.35 0.47 61.06
Enc + MLP Pseudo Loss 15.15 0.64 89.39 27.65 0.53 79.92
RF Uncertainty 17.99 0.20 88.95 50.28 0.10 61.87
CatBoost Uncertainty 14.70 0.33 90.38 25.72 0.14 83.81
100 Binary MLP Uncertainty 20.64 0.49 86.03 46.39 0.30 65.26
Multiclass MLP Uncertainty 14.77 6.44 69.91 35.34 32.64 33.72
Multiclass MLP
+ Binary SVM
Uncertainty 30.45 1.76 74.11 73.47 3.88 31.69
Binary SVM Uncertainty 15.41 0.68 88.38 43.07 0.32 68.33
CADE OOD 23.48 0.96 82.22 58.78 0.70 52.47
multiclass SVM Uncertainty 23.86 0.17 82.18 54.29 0.12 58.26
multiclass GBDT Uncertainty 27.76 0.67 80.15 48.59 0.76 62.58
Enc + MLP Pseudo Loss 13.69 0.44 90.42 27.35 0.41 80.07
RF Uncertainty 15.38 0.20 90.53 48.59 0.09 63.64
CatBoost Uncertainty 11.77 0.26 92.39 26.63 0.13 82.12
200 Binary MLP Uncertainty 19.71 0.39 86.97 42.57 0.34 68.47
Multiclass MLP Uncertainty 14.56 4.26 75.65 39.78 34.76 28.59
Multiclass MLP
+ Binary SVM
Uncertainty 29.46 1.98 74.09 70.32 0.93 39.51
Binary SVM Uncertainty 14.07 0.86 88.47 40.31 0.37 70.24
CADE OOD 21.68 0.67 84.50 51.32 0.78 59.11
multiclass SVM Uncertainty 21.19 0.21 86.90 44.77 0.13 66.55
multiclass GBDT Uncertainty 24.71 0.56 82.71 42.97 0.80 67.28
Enc + MLP Pseudo Loss 9.42 0.48 92.72 27.67 0.39 80.51
RF Uncertainty 13.44 0.19 91.76 49.45 0.90 62.74
CatBoost Uncertainty 10.40 0.24 93.27 22.19 0.13 86.06
400 Binary MLP Uncertainty 16.04 0.40 89.25 36.25 0.34 73.70
Multiclass MLP Uncertainty 15.07 4.15 75.94 34.48 24.44 38.34
Multiclass MLP
+ Binary SVM
Uncertainty 28.85 1.68 75.69 73.94 1.92 33.74
Binary SVM Uncertainty 12.86 0.90 89.02 34.73 0.43 74.12
CADE OOD 20.61 0.59 85.52 49.98 0.94 59.53
multiclass SVM Uncertainty 17.87 0.24 88.88 40.99 0.14 69.61
multiclass GBDT Uncertainty 20.16 0.46 86.24 33.62 0.38 76.82
Enc + MLP Pseudo Loss 7.84 0.50 93.50 21.49 0.31 85.81
RF Uncertainty 11.06 0.20 93.13 47.79 0.90 64.29
CatBoost Uncertainty 9.41 0.23 93.87 26.60 0.10 82.16
TABLE IV: Performance comparison of different models on APIGraph and AndroZoo datasets. The unshaded results are from the original paper, while the shaded results are obtained as part of our benchmarking.
Enc + MLP was evaluated on the two datasets and compared with multiple baseline models. The results showed that it outperforms all baseline models, with a particularly significant improvement on the AndroZoo dataset. The authors have made all research artifacts publicly available, including the datasets used in the experiments and the model-related code. This accessibility allows us to integrate additional testing models into their framework seamlessly, enabling the evaluation of more ML models using the same input and evaluation metrics.
We selected additional ML models from the ones mentioned in Section 4: NB, KNN, RF, and CatBoost (excluding SVM, as it was already included as a baseline in the original work), to expand the set of baseline models. Table IV presents the performance of all tested models, with shaded rows indicating models not included in the original work. If an additional model did not outperform the original baseline models, we omitted its results from the table. As a result, Table IV only includes RF and CatBoost as additional baseline models. The results show that RF achieves performance comparable to the ‘Enc+MLP’ model on the APIGraph dataset but performs poorly on the AndroZoo dataset. However, CatBoost outperforms ‘Enc+MLP’ on both datasets across different labeling budgets, except for the AndroZoo dataset when the labeling budget is set to 400.
7 Benchmarking for CapsGNN
Capsule Graph Neural Networks. The Capsule Graph Neural Network (CapsGNN) approach proposed by Zhang et al. [46] represents an attractive approach to explore in the context of Android malware detection, since it captures information at graph level and enhances graph embeddings obtained with Graph Neural Networks (GNN) [47] by adopting the Capsule Neural Network (CapsNet) [48] ideas. CapsNet helps capture different aspects of the graph, corresponding to different properties, in the final embeddings.
The Capsule Neural Networks (CapsNet) architecture was introduced by Sabour et al. [48] for image feature extraction, although the core concept of a capsule in neural networks was created by Hinton et al. [49]. The CapsNet’s design is based on CNNs, but in addition to detecting features as a CNN does, CapsNet also aims to encode instantiation parameters (such as position, orientation, texture) as part of the detected features. Thus, rather than representing the features using scalar values as in CNNs, the CapsNet uses capsules, which represent a group of neurons, to encode features as vectors. The length of a capsule determines the probability that the corresponding feature is present. The direction of the capsules reflects the instantitation parameters of the feature.
Rather than using a pooling layer to transmit the feature information between layers, CapsNet uses a dynamic routing mechanism to find the best connections between the current and next layer capsules based on agreement. The key innovation of routing by agreement is the capability to capture the spatial relationships between parts and their whole. In our case, the CapsNet is meant to encode precise semantic information from the inter-procedural control flow graphs (ICFGs) of an app.
As GCN, CapsGNN takes as input a graph data structure, while also employing capsules (vectors) instead of scalar values to represent different instantiation parameters of the features. Thus, CapsGNN is a promising DL candidate for Android malware detection, as it can work directly with the ICFG and capture different aspects of the graphs. Figure 1 illustrates the CappsGNN architecture implemented in our Android malware detection system. CapsGNN extracts a primary capsule for each node in an input graph. As opposed to GCN, which extracts node embeddings from the last layer of the network, CapsGNN extracts node embeddings from different layers and represents them as primary capsules (more precisely, the primary capsules are obtained by summing up node embeddings from different layers). As graphs in other application domains, the ICFGs generated from Android apps are diverse, with graph sizes varying widely. As mentioned above, the number of primary capsules depends on the number of nodes in a graph, as each capsule corresponds to a node. Furthermore, the number of graph and class capsules depends on the number of primary capsules. Thus, the node capsules need to be scaled to ensure that the graph capsules from different graphs have the same dimension and are comparable, despite differences in graph sizes. The CapsGNN framework designed by Zhang et al. [46], which we use in this work, implements an attention module between primary capsules and graph capsules. The attention module ensures that: 1) the model identifies the most relevant parts of the graph, and 2) the node capsules are scaled to the same dimension, so that they subsequently lead to comparable graph and class embeddings.
Figure 1: CapsGNN Architecture in Android Malware Detection. N is the number of nodes in one graph, C is the number of node attribute channels, d is the node embedding dimension.
7.1 Benchmarking for CapsGNN
CapsGNN is a DL model designed to capture complex patterns within an app. We include five widely used ML models, as mentioned in Section 4, as potential baselines for comparison with the CapsGNN model.
In addition to traditional ML models, we also experimented with ExcelFormer, a state-of-the-art neural network proposed by Chen et al. [50], which outperforms Gradient Boosted Decision Trees (GBDTs) and existing DL models for tabular data. It addresses key challenges in deep tabular learning through three key innovations: (1) a semi-permeable attention module that reduces the influence of less informative features, (2) data augmentation techniques specifically designed for tabular data, and (3) an attentive feedforward network that enhances model fitting capability. These design choices make ExcelFormer a highly effective solution for diverse tabular datasets. Extensive and rigorous experiments on real-world datasets demonstrate that ExcelFormer consistently outperforms previous models across various tabular prediction tasks. Given that an APK’s structural representations can be effectively converted into tabular data, we selected ExcelFormer to assess its potential for improving Android malware detection performance.
The Android system is an event-based system, and the event-driven control flow can involve various method calls based on the app’s components’ life cycles. Independent method-level control flow graphs (CFGs) or API call sequences could not capture the true invocation order of API calls. In contrast, a ICFG provides a more accurate representation of the actual execution sequence of these. operations. Therefore, we utilize ICFGs as the input representation for our CapsGNN model.
Traditional ML models usually need a predefined feature set to represent an application, and the performance of the classifier highly depends on the features. Section 2 briefly explains some prior works’ approaches to feature selection. The most common features used in prior works are API calls, app permissions, and a few other pieces of information from raw app code. The feature sources are similar but different approaches collect and organize the features in different ways. Among prior works based on traidtional ML, DREBIN [6] has shown high detection preformance on datasets with realistic size and malware to benign app ratio. Recent work [51] shows that DREBIN’s features are still robust for apps as late as 2019. DREBIN extracts static features such as components, intents, API calls, permissions, and network addresses totaling 500,000 features. Roy et al. [28] achieved similar performance as DREBIN by choosing only 471 static features from the original DREBIN’s feature set. To balance the computational cost and performance on large-scale datasets, we use the same features as those in Roy et al.’s work [28], which consist of 471 features including permissions, intent actions, discriminative APIs, obfuscation signatures, and native code signatures. To support future research, the complete feature list will release as open-source with other related code following the acceptance of this work.
7.2 Dataset for CapsGNN
(a) Breakdown of a 6-batch window
(b) 6-batch sliding windows (with 4 training batches)
(c) 7-batch sliding windows (with 5 training batches)
(d) 8-batch sliding windows (with 6 training batches)
Figure 2: Sliding Windows
The dataset plays a crucial role in ML-based malware detection research. To ensure meaningful evaluation, the dataset size needs to reflect the scale of Android apps found in real-world markets. A key challenge in existing datasets is sampling bias [52], which can arise when benign and malicious apps are collected from different sources. To mitigate this issue, data should be gathered from a single market, ensuring consistency in distribution and characteristics. Furthermore, the organization of training and testing subsets should align with real-world deployment scenarios. Specifically, training should be performed on earlier data, while testing should be conducted on later data—a methodology supported by prior research [43]. Additionally, the testing subset must accurately represent the real-world ratio of benign to malicious apps. Some industry experts [43] suggest that the proportion of malware in app markets is relatively low, approximately 6%. Since class imbalance can significantly impact classifier performance, particularly in terms of precision and recall, it is crucial to construct a dataset that mirrors these real-world distributions. By systematically addressing these concerns, we aim to create a more representative and reliable dataset for Android malware detection research.
To assess the impact of app evolution on model performance, we adopt a sliding window approach for constructing training, validation, and test datasets, as illustrated in Figure 2. Specifically, we first partition the data into batches based on app release dates, with each batch containing 5,000 apps—comprising 300 malicious apps and 4,700 benign apps within the same time frame.
We subsequently form windows of consecutive batches, with the first few batches used for training, the batch before the last for validation and the last batch for testing, as shown in Figure 2(a) for a 6-batch window. In each subsequent experiment, we slide the window forward by one batch to form a new dataset for training and evaluation. Our objective is to maximize data availability for both training and testing. A larger training set contributes to more robust models, while an extensive test set ensures reliable performance estimation. Each batch is designed to include 5,000 apps, which we believe is sufficient to provide accurate performance evaluations. To determine the optimal number of training batches for model stability, we experiment with windows containing 4, 5, and 6 training batches, as shown in Figure 2(b), Figure 2(c) and Figure 2(d), respectively.
As illustrated in Figure2(b), our dataset configuration results in 22 overlapping 6-batch windows, with each window consisting of 28,200 benign apps and 1,800 malicious apps—except for the final window, where the availability of malicious apps from late 2021 was limited. Our data partitioning strategy ensures that all experiments are conducted on subsets of consistent size and with a fixed malicious-to-benign ratio. Additionally, this approach enables the observation of temporal trends in Android malware detection.
Figure 3: Top-performing models among all tested models
We will publicly release this dataset upon the publication of the paper.
7.3 Benchmarking Results
Figure 3 illustrates the top-performing ML models when using a window with four training batches. It can be observed that all models exhibit very similar performance. RF, CatBoost, and KNN show nearly identical performance, often overlapping on the graph, while SVM performs slightly lower than the other three models across most windows.
(a) Random forest performance over the sliding windows. The training datasize in a window is varied: 4-batches, 5-batches, and 6-batches.
(b) CapsGNN performance over the sliding windows. The training datasize in a window is varied: 4-batches, 5-batches, and 6-batches.
Figure 4: Main caption for both figures
Figure 5: RF, CatBoost, ExcelFormer and CapsGNN Performance Comparison: Training data consists of 6-batches per window.
Size \ Model
RandomForest
CapsGNN
4 batches 55 13
5 batches 75 17
6 batches 89 20
TABLE V: Model Running Time with Different Training Data Size
How much data is needed for model training? To understand how many training batches need to be included in an experiment window to obtain robust models, we explored windows with different number of training batches as illustrated in Figure 2(b), Figure 2(c) and Figure 2(d) , which show windows with 4, 5 and 6 training batches, respectively.
As can be seen in Figure 4(a) and Figure 4(b) for RF and CapsGNN, a larger training dataset does not always result in increased model performance.
For the traditional ML model, once the size of the training data reaches a certain level, the performance of the models does not improve significantly even when we further increase the size of the training data. In particular, the models trained on 4 batches of data have similar performance as the models trained on 6 batches of data. This can be seen for most windows in Figure 4(a). For CapsGNN, as Figure 4(b) shows, the performance of the model improves when training data size in a window is increased from 4 batches to 6 batches (this is expected given that the CapsGNN model has a large number of parameters that need to be learned). However, the time and resources required to train the model also prohibitively increase (see Table V for time requirements of the two models), which made it impossible for us to experiment with even larger training data for CapsGNN. In contrast, the RF model needs less than two minutes to train regardless of the size of the training subset. Given these results, the remaining experiments are performed using a window size of 8 batches (with 6 batches being used for training).
The results in Figure 5 show that the CapsGNN model has similar performance to the other models for windows from the start of 2018 until the middle of 2019 (i.e., from window 3 through window 14). However, CapsGNN does outperform baseline models in most test batches past late 2019 (in particular, windows 15, 16, 17, and 18). We observe that CapsGNN model has more robust performance than baseline models especially when data evolves (or undergoes significant changes) past late 2019. This is not a surprising result given that the CapsGNN’s input includes much richer program semantics information in the form of ICFG graphs, together with features not present in the ICFG graph itself, including permissions, intent actions and obfuscated/native code signatures. Moreover, CapsGNN dynamically identifies predictive features in the input as part of the learning process, making it easier to keep up with changes in the data. As opposed to that, the input of the RF model includes only manually designed static features that do not change with changes in the data. It is surprising to see that, despite the strictly richer CapsGNN’s input and also the significantly more computational resources required by CapsGNN, CapsGNN fails to produce significant performance improvements over the RF model when the data is relatively stable.
8 Conclusion
In this paper, we systematically evaluate Android malware detection models across four datasets—three publicly available and one systematically collected Google Play-only dataset. We assess the performance of both deep learning (DL) and traditional machine learning (ML) models. After implementing additional models for each proposed DL approach on each dataset, our results show that while DL models achieve strong performance, simpler and computationally efficient ML modelscan achieve comparable or even superior performance. These findings emphasize the importance of extensive benchmarking in Android malware detection research. We encourage future studies to conduct more comprehensive evaluations between advanced DL models and traditional ML models to provide a more accurate assessment of detection capabilities. Additionally, we advocate for releasing datasets used in research to facilitate further research, enabling deeper exploration and new discoveries in this field. We will publicly release our dataset upon the publication of the paper.

### v2

@@
-Benchmarking Android Malware Detection: Rethinking the Role of Traditional and Deep Learning Models
+Benchmarking Android Malware Detection: Traditional vs. Deep Learning Models
 Abstract
@@
 We implement a range of traditional ML models, including Random Forests (RF) and CatBoost, alongside advanced DL models such as Capsule Graph Neural Networks (CapsGNN), BERT-based models, and ExcelFormer-based models.
-Our results reveal that while advanced DL models can achieve strong performance, they are often compared against an insufficient number of traditional ML baselines.
-In many cases, simpler and more computationally efficient ML models achieve comparable or even superior performance.
+Our results reveal that in many cases simpler and more computationally efficient ML models achieve comparable or even superior performance compared with DL models.
 These findings highlight the need for rigorous benchmarking in Android malware detection research.
@@
 Android malware detection, Machine learning, Deep learning, Dataset, Benchmark
-1 Introduction
+I Introduction
 The Android ecosystem maintains the leading position as the largest and most active system among the worldwide mobile computing environment by far.
-Among its ecosystem components, the app markets which consist of Google Play, as well as manufacturer and third-party app stores, play a crucial role in releasing and distributing the apps to end users.
-According to one recent 42matters’s statistics [1], there were 40,325 apps on average released monthly just through Google Play Store between November 2024 and Juanuary 2025.
+According to one recent 42matters’s statistics [1], there were 40,325 apps on average released monthly through Google Play alone between November 2024 and Juanuary 2025.
 The app market’s rapid growth makes it crucial to have an automated and effective app vetting system that scales to the app market sizes.
 An important component of a vetting system is the application of machine learning to triage the apps in the vetting process.
-Traditional machine learning approaches (shorthanded ML hereafter) such as Naive Bayes (NB), k-Nearest Neighbors (k-NN), Support Vector Machines (SVM), Random Forest (RF), Artificial Neural Networks (ANN), Gradient Boosting, etc.
-have been applied to train models on a training dataset, which consists of apps whose maliciousness is known.
-Then the trained models are applied to classify a test dataset, where apps’ maliciousness is unknown.
-Many traditional ML models have shown outstanding performance [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] on datasets from small to market-scale.
+Researchers train a machine learning model on a dataset consistsing of apps whose maliciousness is known.
+Then the trained model is applied to a disjoint test dataset to evaluate the model’s effectiveness.
+Traditional machine learning approaches (shorthanded ML hereafter) such as Naive Bayes (NB), k-Nearest Neighbors (k-NN), Support Vector Machines (SVM), Random Forest (RF), Artificial Neural Networks (ANN), Gradient Boosting, and so on, have shown outstanding performance [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] on datasets from small to market-scale.
 More recently, deep learning approaches (shorthanded DL hereafter) have been studied on Android malware detection [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
 One potential advantage of DL over ML is that it does not require an extensive feature selection and engineering process, although careful consideration still needs to be applied as to how to represent app data provided as input to the DL models.
-Despite the proposed DL approaches for this problem, there exist limited comparisons [12, 21, 27, 24, 26] between ML and the proposed DL models.
-More often than not researchers do not publish the dataset used to compare performances of the various models, making it impossible for other researchers to further benchmark the model against additional alternatives.
-It is encouraging that we are seeing some recent works where researchers publish the datasets used in their study.
-Upon looking through the literature, we identified two works where 1) the researchers compared novel DL models against ML models when applied to Android malware detection, 2) the researchers published the datasets used in their experiments, and 3) the datasets are of substantial sizes.
-This enabled us to conduct additional experiments to benchmark the proposed DL models against additional ML models that were not examined in the original research.
-In addition, we conducted benchmarking of two other DL models against the ML models, on a large dataset that we curated.
-Our experiments show mixed results regarding whether DL models can outperform ML models for Android malware detection.
-The results call for more extensive benchmarking of new deep learning models against traditional machine learning models for Android malware detection.
-2 Related Work
-Machine learning-based Android malware detection systems have been researched for a decade.
-Crowdroid [2] is an earlier ML framework for Android malware detection.
-DroidAPIMiner [3] performs static analysis to extract lightweight features at API level, then uses different classification models such as k-NN, SVM and Decision Trees (DT) to detect malware.
-DREBIN [6] uses similar static analysis techniques as DroidAPIMiner to generate a feature set that contains components, permissions, filtered intents, API calls and network addresses, then uses SVM as a malware classifier.
-Roy et al.
-[28] performed experimentation studies to investigate a number of research questions related to using ML in Android malware detection.
-MaMaDroid [29] abstracts apps’ API calls from 340 packages and 11 families of malware, then applies a Markov chain to form the sequence of abstracted API calls extracted from call graphs as a feature vector, and finally performs classification using Random Forest, k-NN and SVM algorithms.
-APIChecker [11] uses dynamic analysis to collect API features, applies nine machine learning algorithms for Android malware detection and evaluates their performance on the T-Market data with about 500K apps.
-The results show that RF yields the best overall performance among the selected models.
-DroidAPIMiner [4], DREBIN [6], DroidMat [30], and MaMaDroid [29] also compared their performance with other traditional ML classifiers and anti-virus scanners.
-MaMaDroid uses four different ML algorithms and the detection results obtained with RF have better performance in general.
-Recent years have seen an increase in deep learning related Android malware detection research.
-The deep neural network models used include deep belief networks (DBN) [12, 13], convolutional neural networks (CNN) [15, 16, 24, 31], pre-trained BERT models [26], recurrent neural networks (mainly long short-term memory networks, or LSTM) [32, 19], and multilayer perceptrons (MLP) with continuous learning schemes [25], among others.
-Droid-Sec [12] trained and tested a DBN model based on 500 apps (250 malware).
-Experimental results showed that the DL model has significant accuracy improvement as compared to traditional ML models, such as SVM, NB, Logistic Regression, and MLP.
-Deep4MalDroid [14] evaluated a Stacked AutoEncoders (SAEs) model based on 3000 apps (50% of which are malware).
-Results showed that the proposed DL model has higher accuracy than the selected ML models, specifically, SVM, ANN, NB, and DT.
-DeepFlow [33] proposed a DBN model, and conducted experiments based on 3000 benign apps and 8000 malicious apps.
-The comparison between DL and traditional ML models showed that DBN and SVM outperform other selected ML models in balancing the recall and precision.
-DexRay [24] converts the raw bytecode of DEX files in the APK as one stream vector, then converts the vector to a grey-scale image which is provided as input to a CNN architecture.
-DexRay evaluated the performance of the CNN model based on 61,809 malicious and 134,134 benign apps, then compared its detection results with DREBIN’s results and other image-based malware detection approaches.
-DexRay yields almost the same performance metrics as DREBIN; however, its performance can be significantly affected if obfuscated apps are absent in the training data but present in the test data.
-DetectBERT [26] leverages a pre-trained BERT-like model with correlated Multiple Instance Learning (c-MIL) to process Smali code from APKs, enriching class-level features and aggregating them into an app-level representation for effective Android malware detection.
-It employs the DexRay dataset as its benchmark baseline for evaluation.
-DeepRefiner [17] is a two-layer detection system, with the first detection layer based on MLP, and the second layer based on an LSTM network.
-The experimental data include 47,525 benign apps and 62,915 malicious apps.
-The detection performance, based on accuracy, TPR and FPR, is approximately 10% higher than that of a fine-turned SVM model.
-DANdroid [34] uses the DREBIN dataset to evaluate its proposed DL discriminative adversarial network, and the performance is very close to that of DREBIN.
-DL-Droid [23] generates 420 dynamic stateful features such as application attributes, actions and permissions by running apps on real devices, then feeding them into a multilayer perceptron classifier.
-It uses weighted F measure (w-FM) as main evaluation metric and compares it against that of seven selected ML algorithms.
-DL-Droid reached the highest w-FM value of 98.82% (the best ML model RF reached w-FM value of 97.1%).
-AMD-CNN [31] is a visualization-based Android malware detection tool.
-It extracts the features from Android manifest file and converts them to RGC images, then feeds the images to a CNN network.
-It was evaluated based on 960 benign and 960 malicious apps.
-[35] identified five factors that influence the training and performance of ML-based Android malware detectors.
-The authors conducted a fair comparison of eight detectors on a mixed-market dataset, accounting for these factors.
-The detectors include seven traditional ML models (across four types) and one DBN with two hidden layers of 150 neurons each.
+Despite various DL approaches proposed for Android malware detection, there is limited comparison between DL and traditional ML models [12, 21, 27, 24, 26].
+Evaluating DL models with only a few baselines often fails to reflect true performance – a pitfall termed “inappropriate baseline” [28] previously noted in areas like vulnerability discovery and network intrusion detection.
+We observed similar shortcomings in Android malware detection research.
+Many studies do not release their datasets, making it difficult to benchmark against a broader range of alternatives.
+Encouragingly, some recent works have made their datasets publicly available.
+Through our literature review, we identified two studies that 1) compared DL and ML models, 2) published their datasets, and 3) used datasets of substantial sizes.
+This allowed us to further benchmark those DL models against additional ML models.
+We also evaluated two other DL models against ML models using a large dataset we curated.
+Our experiments show mixed results regarding whether DL consistently outperforms ML models, underscoring the need for more comprehensive benchmarking in Android malware detection.
+II Related Work
+Machine learning-based Android malware detection has been studied for over a decade.
+Early approaches like Crowdroid [2] used k-means clustering to separate benign apps from malware, DroidAPIMiner [3] extracted static API-level features and applied classifiers such as k-NN, SVM, and Decision Trees (DT).
+DREBIN [6] represented apps using a rich static feature set and classified them using SVM.
+MaMaDroid [29] abstracted API calls into Markov chains and used classifiers like RF and SVM.
+APIChecker [11] collected dynamic API features and evaluated nine ML algorithms, with RF performing best on a dataset of approximately 500K apps.
+Other systems, including DroidMat [30], APIGraph [31], and MalScan [32], employed lightweight static or graph-based features with classifiers like k-NN and RF, consistently demonstrating strong detection performance.
+Recent research has increasingly focused on deep learning (DL) for Android malware detection, leveraging models such as deep belief networks (DBN) [12, 13], convolutional neural networks (CNN) [15, 16, 24, 33], long short-term memory (LSTM) networks [34, 19], pre-trained BERT models [26], and multilayer perceptrons (MLP) [25].
+Droid-Sec [12] used DBN and showed improved accuracy over traditional ML models.
+Deep4MalDroid [14] applied stacked autoencoders (SAEs) and outperformed classifiers like SVM and NB.
+DeepFlow [35] showed DBN and SVM achieved better balance between recall and precision than other ML models.
+DexRay [24] transformed DEX bytecode into grayscale images for CNN-based detection, achieving performance comparable to DREBIN.
+DetectBERT [26] used BERT [36] with Multiple Instance Learning to process Smali code, improving class-level representation.
+DeepRefiner [17] combined MLP and LSTM in a two-layer system, achieving  10% higher accuracy than SVM.
+DANdroid [37] evaluated a discriminative adversarial network on the DREBIN dataset, showing similar results.
+DL-Droid [23] extracted 420 dynamic features and used MLP, outperforming seven ML baselines.
+AMD-CNN [33] visualized Android manifest features as RGC images for CNN input.
 Chen et al.
-[25] proposed a continuous learning approach using MLP to enhance the adaptability of ML-based Android malware detectors to concept drift [36].
-Their work includes a comprehensive comparison with major baseline models, with both input data and model-related code available as open source.
-The evaluation was conducted using two established datasets, APIGRAPH [37] and AndroZoo [38].
-Our study complements these prior works by extensive benchmarking the proposed DL models against additional standard ML models that were not examined by the original researchers.
-Our results lead to reassessment of DL’s advantage against ML in Android malware detection, and highlights the need for extensive benchmarking for this problem domain.
-3 Benchmarking Methodology
+[25] proposed a continuous learning MLP model for handling concept drift [38], evaluated on APIGraph [31] and AndroZoo [39] with major baseline models.
+III Benchmarking Methodology
 Benchmarking across different models is a standard approach for evaluating and comparing their performance, as well as identifying the strengths and weaknesses of various approaches.
@@
 For this reason, it is utterly important that researchers make the dataset upon which their models are evaluated available.
-This will enable other researchers to evaluate additional models that had not been considered in the original researcher.
+This will enable other researchers to evaluate additional models that had not been considered in the original research.
 More extensive evaluations can provide deeper insights and help understand the true strengths of the various models for Android malware detection.
-One of the key challenges for such extensive evaluation is that not all works publish the datasets used in their study.
-Fortunately, we have identified recent studies, such as Chen et al.’s Enc+MLP [25] and Sun et al.’s DetectBERT [26], that have made their datasets publicly available.
-This allowed us conduct more extensive benchmarking of the models proposed in these two works.
-Additionally, we collected a Google Play-only dataset and used two DL approaches, Capsule Graph Neural Networks (CapsGNN) and ExcelFormer, which have not been explored before for Android malware detection.
-Each of the three works employs different datasets to evaluate their models, and their approaches to app representation also vary.
 Our primary objective is not to assert the superiority of any specific model but to demonstrate how comprehensive benchmarking enhances the understanding of model performance.
@@
 In fact to understand how choices in datasets construction impact models’ measured performance is a non-trivial research question in its own right and beyond the scope of this paper.
-In the following sections, we first discuss traditional machine learning models considered and then explore additional models that were not included in the two prior studies.
-4 Traditional Machine Learning Models Considered
-Widely explored ML models in Android malware detection include Naive Bayes (NB) [39, 40], support vector machines (SVM) [4, 6], k-nearest neighbors (k-NN) [30, 40], and random forest (RF) [29].
+III-A Traditional Machine Learning Models Considered
+Widely explored ML models in Android malware detection include Naive Bayes (NB) [40, 41], support vector machines (SVM) [4, 6], k-nearest neighbors (k-NN) [30, 41, 32], and random forest (RF) [29, 32].
 We consider these four widely used models as potential baselines to compare with the proposed DL models.
@@
 Thus we also include it as one of the ML models in our benchmarking.
-NB: This is a simple but effective classifier. Furthermore, it is the fastest classifier we evaluated. It can get the classification result quickly even with half a million apps.
-SVM: Widely used classifier in Android malware classification.
-It requires more memory and computation time during the training stage, which poses a significant challenge for large datasets.
-We set the kernel to be linear, and use a grid search with values 0.001, 0.1, 1.0, and 10 to decide the optimal regularization parameter to train the SVM model.
-k-NN: Another simple and robust traditional classifier.
-The value of the hyper-parameter $k$ has a direct impact on the performance.
-We use a grid search with values between 3 and 15 to determine the optimal $k$ value to train the k-NN model.
-RF: A popular ensemble-type model based on decision trees.
-It has shown promising performance in Android malware detection [29] and has the capability to be applied to market-scale datasets [11].
-We perform a grid search to find the optimal number of trees in the forest.
-The estimator value is between 100 and 300.
-CatBoost: Another ensemble-based model built on decision trees.
-While it is relatively new compared to the other ML models, recent studies across various domains have demonstrated its effectiveness in classification tasks [41].
-CatBoost is sensitive to hyperparameters, and our experimental results are based on the following setup: 1000 iterations, a learning rate of 0.1, a depth of 10, and an L2 regularization parameter (l2_leaf_reg) of 5.
-5 Additional Benchmarking for DetectBERT
+III-B Deep Learning Models Considered
+One of the key challenges in applying recent DL methods to Android malware detection is the lack of publicly available datasets in their studies.
+Fortunately, we have identified recent studies, such as Chen et al.’s Enc+MLP [25] and Sun et al.’s DetectBERT [26], that have made their datasets publicly available.
+This allowed us to conduct more extensive benchmarking of the models proposed in these two works.
+Additionally, we collected a Google Play-only dataset and used two DL approaches, Capsule Graph Neural Networks (CapsGNN) and ExcelFormer, which have not been explored before for Android malware detection.
+Each of the three works employs different datasets to evaluate their models, and their approaches to app representation also vary.
+IV Case Studies
+IV-A Case Study 1: DetectBERT
 DetectBERT is a recent approach proposed by Sun et al.
@@
 This approach enhances class-level features and aggregates them into an app-level representation, improving the effectiveness of Android malware detection.
-5.1 Benchmarking for DetectBERT
+IV-A1 Benchmarking for DetectBERT
 DetectBERT was evaluated against two state-of-the-art Android malware detection models, Drebin [6] and DexRAY [24].
@@
 The results indicate that DetectBERT slightly outperforms both baseline models.
-5.2 Dataset for DetectBERT
+IV-A2 Dataset for DetectBERT
 DetectBERT Benchmark Dataset.
@@
 The labeling process is based on VirusTotal reports: apps that are not flagged by any antivirus engines are considered benign, while those detected as malicious by more than two antivirus engines are labeled as malware.
-5.3 Additional Benchmarking
-Model Accuracy Precision Recall F1 Score
-Drebin 0.97 0.97 0.94 0.96
-DexRay 0.97 0.97 0.95 0.96
-DetectBERT 0.97 0.98 0.95 0.97
-NB 0.85 0.87 0.85 0.86
-RF 0.97 0.97 0.97 0.97
-KNN 0.97 0.97 0.97 0.97
-SVM 0.96 0.96 0.96 0.96
-CatBoost 0.98 0.98 0.98 0.98
-TABLE I: Performance comparison of various models on DetectBERT benchmark dataset. The unshaded are from the original paper, while the shaded results are from our benchmarking.
+IV-A3 Additional Benchmarking
 Sun et al.
@@
 These ML models also require significantly less compute time than DetectBERT.
+Model Accuracy Precision Recall F1 Score
+Drebin 0.97 0.97 0.94 0.96
+DexRay 0.97 0.97 0.95 0.96
+DetectBERT 0.97 0.98 0.95 0.97
+NB 0.85 0.87 0.85 0.86
+RF 0.97 0.97 0.97 0.97
+KNN 0.97 0.97 0.97 0.97
+SVM 0.96 0.96 0.96 0.96
+CatBoost 0.98 0.98 0.98 0.98
+TABLE I: Model performance on the DetectBERT benchmark dataset. Unshaded results are from the original paper; shaded results are from our benchmarking.
 In the paper the authors also conducted experiments on data sets that adhere to temporal consistency.
 However the published dataset does not include the train/test data split for this experiment, and as a result we were not able to compare additional ML models’ results against the results published in the paper.
-6 Additional Benchmarking for Enc+MLP
+IV-B Case Study 2: Enc+MLP
 Enc + MLP is proposed by Chen et al.
@@
 This approach improves upon existing active learning baselines for Android malware detection while minimizing the need for manual labeling, as demonstrated through evaluations on two publicly available datasets.
-6.1 Benchmarking for Enc+MLP
+IV-B1 Benchmarking for Enc+MLP
 Enc+MLP uses multiple baseline models and explores uncertainty sampling for both binary and multiclass classifiers.
@@
 All baseline models employ a cold-start approach (retraining from scratch on the updated training data) for active learning.
-6.2 Datasets for Enc+MLP
+IV-B2 Datasets for Enc+MLP
 Enc+MLP was evaluated on two datasets: APIGraph and AndroZoo.
@@
 The malicious apps are sourced from VirusTotal [43], VirusShare [45], and the AMD dataset [46], while the benign apps are obtained from AndroZoo [39].
-Table III provides a detailed breakdown of this dataset.
-AndroZoo Dataset.
-Collected by Chen et al.
-[25] from AndroZoo [38], this dataset includes Android apps from 2019 to 2021.The apps are also ordered based on their appearance timestamps.
-Malicious apps were randomly selected based on VirusTotal reports, where at least 15 antivirus engines flagged them as malware.
-Benign apps were randomly chosen from those with 0 positive detections by any antivirus engine.
-The dataset maintains a similar malware/benign app ratio, with 90% benign apps and 10% malicious apps each year.
-Table III provides a detailed breakdown of the collected data for each year.
-6.3 Additional Benchmarking
-Monthly Sample Budget Model Architecture Sample Selector APIGraph Dataset Average Performance (%) AndroZoo Dataset Average Performance (%)
+Table II provides a detailed breakdown of this dataset.
+Monthly
+Sample
+Budget
+Model
+Architecture
+Sample
+Selector
+APIGraph Dataset
+Average Performance (%)
+AndroZoo Dataset
+Average Performance (%)
 FNR FPR F1 FNR FPR F1
@@
 The unshaded results are from the original paper, while the shaded results are obtained as part of our benchmarking.
+AndroZoo Dataset.
+Collected by Chen et al.
+[25] from AndroZoo [39], this dataset includes Android apps from 2019 to 2021.The apps are also ordered based on their appearance timestamps.
+Malicious apps were randomly selected based on VirusTotal reports, where at least 15 antivirus engines flagged them as malware.
+Benign apps were randomly chosen from those with 0 positive detections by any antivirus engine.
+The dataset maintains a similar malware/benign app ratio, with 90% benign apps and 10% malicious apps each year.
+Table III provides a detailed breakdown of the collected data for each year.
+IV-B3 Additional Benchmarking
 Enc + MLP was evaluated on the two datasets and compared with multiple baseline models.
@@
 This accessibility allows us to integrate additional testing models into their framework seamlessly, enabling the evaluation of more ML models using the same input and evaluation metrics.
-We selected additional ML models from the ones mentioned in Section 4: NB, KNN, RF, and CatBoost (excluding SVM, as it was already included as a baseline in the original work), to expand the set of baseline models.
+We selected additional ML models from the ones mentioned in Section III-A: NB, KNN, RF, and CatBoost (excluding SVM, as it was already included as a baseline in the original work), to expand the set of baseline models.
 Table IV presents the performance of all tested models, with shaded rows indicating models not included in the original work.
@@
 As a result, Table IV only includes RF and CatBoost as additional baseline models.
-The results show that RF achieves performance comparable to the ‘Enc+MLP’ model on the APIGraph dataset but performs poorly on the AndroZoo dataset.
-However, CatBoost outperforms ‘Enc+MLP’ on both datasets across different labeling budgets, except for the AndroZoo dataset when the labeling budget is set to 400.
-7 Benchmarking for CapsGNN
+The results show that RF achieves performance comparable to the Enc+MLP model on the APIGraph dataset but performs poorly on the AndroZoo dataset.
+However, CatBoost outperforms Enc+MLP on both datasets across different labeling budgets, except for the AndroZoo dataset when the labeling budget is set to 400.
+IV-C Case Study 3: CapsGNN
 Capsule Graph Neural Networks.
@@
 Figure 1: CapsGNN Architecture in Android Malware Detection. N is the number of nodes in one graph, C is the number of node attribute channels, d is the node embedding dimension.
-7.1 Benchmarking for CapsGNN
+IV-C1 Benchmarking for CapsGNN
 CapsGNN is a DL model designed to capture complex patterns within an app.
-We include five widely used ML models, as mentioned in Section 4, as potential baselines for comparison with the CapsGNN model.
+We include five widely used ML models, as mentioned in Section III-A, as potential baselines for comparison with the CapsGNN model.
 In addition to traditional ML models, we also experimented with ExcelFormer, a state-of-the-art neural network proposed by Chen et al.
@@
 Traditional ML models usually need a predefined feature set to represent an application, and the performance of the classifier highly depends on the features.
-Section 2 briefly explains some prior works’ approaches to feature selection.
+Section II briefly explains some prior works’ approaches to feature selection.
 The most common features used in prior works are API calls, app permissions, and a few other pieces of information from raw app code.
@@
 To support future research, the complete feature list will release as open-source with other related code following the acceptance of this work.
-7.2 Dataset for CapsGNN
+IV-C2 Dataset for CapsGNN
 (a) Breakdown of a 6-batch window
@@
 Specifically, we first partition the data into batches based on app release dates, with each batch containing 5,000 apps—comprising 300 malicious apps and 4,700 benign apps within the same time frame.
+Figure 3: Top-performing models among all tested models
+(a) Random forest performance over the sliding windows. The training datasize in a window is varied: 4-batches, 5-batches, and 6-batches.
+(b) CapsGNN performance over the sliding windows. The training datasize in a window is varied: 4-batches, 5-batches, and 6-batches.
+Figure 4: RF and CapsGNN performance over the sliding windows
 We subsequently form windows of consecutive batches, with the first few batches used for training, the batch before the last for validation and the last batch for testing, as shown in Figure 2(a) for a 6-batch window.
@@
 Additionally, this approach enables the observation of temporal trends in Android malware detection.
-Figure 3: Top-performing models among all tested models
 We will publicly release this dataset upon the publication of the paper.
-7.3 Benchmarking Results
+IV-C3 Benchmarking Results
 Figure 3 illustrates the top-performing ML models when using a window with four training batches.
@@
 RF, CatBoost, and KNN show nearly identical performance, often overlapping on the graph, while SVM performs slightly lower than the other three models across most windows.
-(a) Random forest performance over the sliding windows. The training datasize in a window is varied: 4-batches, 5-batches, and 6-batches.
-(b) CapsGNN performance over the sliding windows. The training datasize in a window is varied: 4-batches, 5-batches, and 6-batches.
-Figure 4: Main caption for both figures
 Figure 5: RF, CatBoost, ExcelFormer and CapsGNN Performance Comparison: Training data consists of 6-batches per window.
@@
 It is surprising to see that, despite the strictly richer CapsGNN’s input and also the significantly more computational resources required by CapsGNN, CapsGNN fails to produce significant performance improvements over the RF model when the data is relatively stable.
-8 Conclusion
+V Conclusion
 In this paper, we systematically evaluate Android malware detection models across four datasets—three publicly available and one systematically collected Google Play-only dataset.